OpenStudy (anonymous):

graph the ellipses, state the focal points of each ellipse. equation: (x-5)^2/25 + (y-2)^2/4=1

OpenStudy (anonymous):

the center: (5,2) c^2= 5^2 -2^2=√21 foci at:(5 ∓ √21 , 2) The lenghth of the major axis is : 5(2) = 10 The lenghth of the minor axis is : 2(2) = 4 x^2 greater denominator: the major axis is horizontal.